OpenStudy (aundiewear):

De Moivre's Theorem ((2)(sqrt3) +2i) ^-5

OpenStudy (irishboy123):

if i read this correctly, you have \(((2)(\sqrt{3}) +2i) ^{-5} \) \( \to 2^{-5} (\sqrt{3} +i) ^{-5} \) \( \to 2^{-5} 2^{-5} (\dfrac{\sqrt{3}}{2} +\dfrac{1}{2}i) ^{-5} \) then \((\cos x + i \sin x)^n = \cos (nx) + i \sin (nx)\) or \(\left( e^{ix} \right)^n = e^{inx} \) if you prefer with \(x = \pi / 6\)